----------------------------------------------------------------------------- Mauna Loa Solar Observatory Observer's Log ----------------------------------------------------------------------------- Tue Oct 27 16:41:30 GMT 2026 Year:26 Doy:300 Observer: mcotter Weather COMMENT BY mcotter: Tue Oct 27 16:49:29 GMT 2026 MLO Current Weather Timestamp (UTC): 2026-10-27 16:38:00 Temperature : 48.27 deg F Dew Point : 41.14 deg F Pressure : 680.28 mbar Wind Speed : 17.9 mph Wind Direction : 208.1 deg Skies: The sky appears hazy and there are lenticular shaped clouds scattered across the sky in many areas, including the eastern horizon. Additionally, there are bands of Cirrus and Cirrostratus clouds scattered across the sky. The wind is blowing gustily from the south-southwest. ___end___ GENERAL COMMENT BY mcotter: Tue Oct 27 17:03:28 GMT 2026 Dome open. The dome are windows remain closed at this time due to the strong gusty wind conditions. ___end___ ** KCOR PM mcotter ** : Tue Oct 27 17:03:37 GMT 2026 Blew off Kcor O1 ___end___ ** UCOMP PM mcotter ** : Tue Oct 27 17:03:45 GMT 2026 Blew off UCoMP O1 ___end___ GENERAL COMMENT BY mcotter: Tue Oct 27 17:08:49 GMT 2026 Due to poor sky conditions I don't believe observations are going to take place. However, just in case the sky does clear a bit later in the morning I have opened the dome and started the spar tracking to be ready for any good sky conditions that may arise. ___end___ Weather COMMENT BY mcotter: Tue Oct 27 17:30:18 GMT 2026 MLSO Current Weather Timestamp (UTC): 2026-10-27 17:23:00 Temperature : 52.54 deg F Dew Point : 41.06 deg F Pressure : 680.63 mbar Wind Speed : 15.66 mph Wind Direction : 200.6 deg Skies: Unfortunately sky conditions are worsening, particularly to the east. Any hopes of the Sun possibly rising above the cloud formations on the eastern horizon don't look good at this time. ___end___ GENERAL COMMENT BY berkey: Tue Oct 27 23:14:08 GMT 2026 The network currently doesn't allow kodiak to send email to google groups. Logger code updated to send the log file to kodiak but email it from Boulder. ___end___ KCOR COMMENT BY berkey: Tue Oct 27 23:20:42 GMT 2026 A missing step was found in the kcor data transfer script. The data was not getting copied to the location that the pipeline expects it. The script has been updated to put a copy in the raw folder. At this point the kcor watch dog transfer program should not be run; as data transfer is handle by Kodiak via the z drive. ___end___ KCOR COMMENT BY KCOR: Wed Oct 28 00:31:26 GMT 2026 KCOR file transfer report KCor created 39 .bin files 39 .fts files WARNING 2 fts file(s) found locally that was not transfered dawn ACTION must be taken to copy: 20261027_201748_kcor.fts 20261027_201803_kcor.fts ___end___ ** GENERAL PM berkey ** : Wed Oct 28 01:11:46 GMT 2026 The Simco Ionizer air gun that is used to blow off the instrument O1's started acting erratically and the triggered airflow stays on even when the trigger is released. The trigger on the gun was cleaned and lubricated. The Simco main unit was opened and inspected. It appears that the trigger controlled solenoid valve is sticking. When the unit is powered off the solenoid does reset and close. There were part numbers and a manufacturer information tag on the solenoid. I looked up the sales department of Solenoid Solutions, Inc. and sent them an email with photos of the solenoid and requested a quote for a new solenoid. The Simco Ionizer air gun was returned to service, but unfortunately when using it will have to be powered off each time the trigger is released in order for the solenoid to close. Also, the descant dryer module that is mounted on the small air compressor and is inline with the Simco Ionizer gun was replaced with a new unit. Additionally, the water separator bowl had small water droplets showing in the bowl. The bowl was removed from the compressor, thoroughly cleaned and reinstalled. ___end___ BY berkey: Wed Oct 28 01:42:00 GMT 2026 End Of Day: Due to very poor sky conditions no observations were taken today. ___end___ Logger COMMENT BY berkey: Wed Oct 28 01:42:02 GMT 2026 Log closed ONSITE STAFF: berkey, mcotter ___end___